# Perturbation-based Effect Measures for Compositional Data

### v3

@@ Abstract
Existing effect measures for compositional features are inadequate for many modern applications for two reasons. First, modern datasets with compositional covariates, for example in microbiome research, display traits such as high-dimensionality and sparsity that can be poorly modelled with traditional parametric approaches. Second, assessing – in an unbiased way – how summary statistics of a composition (e.g., racial diversity) affect a response variable is not straightforward. In this work, we propose a framework based on hypothetical data perturbations that addresses both issues. Unlike many existing effect measures for compositional features, we do not define our effects based on a parametric model or a transformation of the data. Instead, we use perturbations to define interpretable statistical functionals on the compositions themselves, which we call average perturbation effects. These effects naturally account for confounding that biases frequently used marginal dependence analyses. We show how average perturbation effects can be estimated efficiently by deriving a perturbation-dependent reparametrization and applying semiparametric estimation techniques. We analyze the proposed estimators empirically on simulated and semi-synthetic data and demonstrate advantages over existing techniques on data from New York schools and microbiome data. For all proposed estimators, we provide confidence intervals with uniform asymptotic coverage guarantees.
@@ 1 Introduction
Modelling the relationship between a real-valued response $Y\in\mathbb{R}$ and a composition $Z=(Z^{1},\dots,Z^{d})\in\Delta^{d-1}$ has a long history and has, for example, been considered in the literature on ‘experimental design with mixtures’ (Claringbold, 1955; Cornell, 2002) and ’log-ratio based compositional data analysis’ (Aitchison, 1982). We briefly summarize the developments of these fields in Section 1.1. A key aspect of this research is that it attempts to achieve two separate goals simultaneously: (i) ensuring that the model parameters are interpretable and (ii) providing parametric models that fit to realistic data-generating processes with compositional covariates. More recently, in the context of high-dimensional and sparse data, alternative approaches (Knight et al., 2018; Cammarota et al., 2020; Huang et al., 2023) to solving the second goal have been proposed that apply flexible machine learning methods to model the relationship between $Y$ and $Z$ nonparametrically. However, such approaches are at odds with the first goal of providing interpretable effect measures and cannot be combined with the above-mentioned approaches which heavily rely on parametric models to define such effects.
@@
Our main contribution is to provide a unified framework for defining interpretable effect measures for compositional features which we call average perturbation effects that can be combined with flexible nonparametric estimation procedures. Special cases of these average perturbation effects have interpretations that are similar to some model-based individual feature effect measures, e.g., Cox-directions (Cox, 1971) and log-contrast coefficients (Aitchison and Bacon-Shone, 1984). Average perturbation effects, however, are defined without reference to a specific regression model and therefore do not rely on a correctly specified parametric model. Importantly, this allows us to construct more flexible target quantities, two of which we believe are particularly relevant for applications. First, effects based on binary perturbations are useful when summarizing the effect of a discrete change, e.g. the effect of setting $Z^{j}$ to 00 while rescaling $Z^{-j}:=(Z^{1},\dots,Z^{j-1},Z^{j+1},\dots,Z^{d})$ to the simplex. In Section 4.2 we show empirically that the compositional features with significant effects as measured by these binary perturbations coincide with those deemed important by the standard log-contrast regression procedures used in the microbiome literature. Second, effects based on directional perturbations summarize continuous changes in specific directions. This can involve changing individual components of $Z$, but also allows for more complex changes that are not tied to a specific coordinate of $Z$. For example, we can define the effect of ‘diversifying’ $Z$, i.e. moving closer to the center of $\Delta^{d-1}$. Such diversifying effects have been considered previously (Antonio et al., 2004) but are usually defined as the effect of marginally changing a specific measure of diversity. However, as the effect of diversifying can be confounded by other features of $Z$ such a marginal analysis can be misleading and due to the simplex constraint, there is no immediate way to correct for this confounding when estimating these effects (see Example 2 and Section 4.1). In contrast, our perturbation-based framework can control for the remaining variation in $Z$ leading to unconfounded effect estimates with valid asymptotic inference.
@@
Our proposal is intimately related to the growing literature on targeted learning and semiparametric inference (Chernozhukov et al., 2018; Berk et al., 2021; Vansteelandt and Dukes, 2022). An important aspect of our work is, however, to clarify the need for tailored target estimands in the presence of compositional structure in the covariates. Applying a nonparametric effect measure that is not aware of the sum-to-one constraint of the compositional vector can be just as misleading as naively estimating a non-compositional effect based on a parametric model (see Example 1 and Section 4.2). Beyond defining average perturbation effects, we also provide several algorithm-agnostic procedures to estimate the average perturbation effects. These procedures are essentially well-known semiparametric learning algorithms applied after appropriately transforming the compositional covariates. While these algorithms require us to estimate nuisance functions relating to the distribution of $(Y,Z)$, which are particularly delicate for compositional covariates, they only require these estimates to be accurate and not interpretable. This allows us to apply arbitrary regression methods as long as they are sufficiently predictive.
@@
Average perturbation effects are defined by comparing the change in the expected response of $Y$ under a hypothetical perturbation of the composition $Z$. To motivate this, consider unconstrained covariates $X\in\mathbb{R}^{d}$ instead of the composition $Z\in\Delta^{d-1}$. Suppose that there exists $\beta=(\beta^{1},\dots,\beta^{d})\in\mathbb{R}^{d}$ such that for all $x=(x^{1},\dots,x^{d})\in\mathbb{R}^{d}$, $f(x):=\mathbb{E}[Y\,|\,X=x]=\sum_{j=1}^{d}\beta^{j}x^{j}$, that is, the conditional expectation of $Y$ given $X=x$ is a linear function. Then, one interpretation of the coefficient $\beta^{j}$ is that it corresponds to the expected rate of change of the response in the $j$th coordinate, or more formally,
$\partial_{x_{j}}f(x^{1},\dots,x^{d})=\beta^{j}.$
We can rephrase this interpretation by considering a perturbation $\psi:\mathbb{R}^{d}\times[0,1)\rightarrow\mathbb{R}^{d}$ that additively increases the $j$th coordinate, that is, $\psi(x,\gamma)=(x^{1},\dots,x^{j}+\gamma,\dots,x^{d})$. Then, for any $x\in\mathbb{R}^{d}$, we have
@@
Returning now to the composition $Z$ and redefining for $z=(z^{1},\dots,z^{d})\in\Delta^{d-1}$, $f(z):=\mathbb{E}[Y\,|\,Z=z]$, we see that neither of the above expressions are well-defined. The quantity $f(\psi(z,\gamma))$ is meaningless for any non-zero $\gamma\in\mathbb{R}$, since $f$ is only defined on $\Delta^{d-1}$. We can however use similar ideas by defining tailored perturbations on $\Delta^{d-1}$ to obtain interpretable effect measures. The following two examples provide toy applications that illustrate the pitfall and short-comings of existing approaches.
Example 1 (Effects of presence or absence of microbes).
Consider the problem of determining whether the presence or absence of certain microbes in the human gut is affecting a disease. Microbial abundances are commonly measured using genetic sequencing (e.g., targeted amplicon or shotgun metagenomic sequencing) which in most cases only measures the composition and not the absolute abundances. Assume we have measurements from a cohort of patients, where for each patient the relative abundance of all microbes in their gut $Z\in\Delta^{d-1}$ and a disease indicator $Y\in\{0,1\}$ ($1$ encoding disease) has been recorded. While estimating causal effects in such a setting is often too ambitious, one could hope to screen for relevant microbes that could then be considered for follow-up randomized studies. One possible approach (e.g., Hu et al., 2021) would be to consider a fixed microbe $j$, define the treatment variable $T:=\mathds{1}_{\{Z^{j}=0\}}$ corresponding to the presence or absence of microbe $j$ and then estimate a marginal treatment effect given by
that is, the absence of $Z^{1}$ increases while the absence of $Z^{2}$ reduces the probability of having the disease. In a numerical simulation of this generative model with dependence between $Z^{1}$ and $Z^{2}$, the effect of $Z^{1}$ is instead estimated to decrease the probability of disease as shown in Table 1 (details are provided in Section S4.1 of the supplementary material). If $Z\in\mathbb{R}^{d}$ was unconstrained, we could adjust for the dependence in the remaining coordinates by estimating a conditional treatment effect such as
and marginalizing it to produce a single effect measure $\mathbb{E}[h(Z^{-j})]$. [^2] In the literature on causal inference (and under appropriate causal assumptions), $h$ corresponds to the conditional average treatment effect (CATE). Estimating this quantity is a well-studied problem, e.g. using the AIPW estimator (Robins et al., 1994) or using a partially linear model (Chernozhukov et al., 2018). Unfortunately, neither of the estimators is well defined when $Z\in\Delta^{d-1}$ as it is not possible to change $T$ and $Z^{-j}$ independently of each other and, in fact, for all $z^{-j}$, we have $h(z^{-j})=0$ as $Y$ is conditionally independent of $T$ given $Z^{-j}$ due to the sum-to-one constraint which makes $T$ fully determined by $Z^{-j}$. In the Table 1, we provide the results from applying our proposed marginalized conditional treatment effect the compositional knockout effect to $Z^{1}$ $(\mathrm{CKE}^{1})$ which we introduce in Section 2.2.
Example 1
method estimate 95%-CI
Example 2
method estimate 95%-CI
Table 1: Numerical illustration of Examples 1 and 2. The first row in each table corresponds to the marginal treatment effect (MTE or OLS of $Y$ on $-G(Z)$) estimate based on a correctly specified linear model. The second row correspond to estimates of two of the target parameters introduced in Section 2.2 with random forest (rf) as nuisance function estimators. Bold red numbers indicate a significant negative effect at a $5\%$ level, while bold green indicates a significant positive effect. Left: The estimated MTE is negative, even though the true effect is positive. The compositional knock-out effect ($\mathrm{CKE}^{1}$), which adjust for the simplex structure, results in the correct sign. Right: The estimated marginal effect of an increase in diversity (measured by the negative Gini coefficient, i.e., $-G(Z)$) is negative even though the true effect is equal to $1$ by construction. The compositional diversity effect $\mathrm{CDI}_{\mathrm{Gini}}$ correctly recovers the effect. Details on both experiments are provided in Section S4.1 of the supplementary material.
Example 2 (Effects of increased diversity).
Consider the problem of determining the effect of an increase in diversity of $Z$ on a response $Y$. Such a question is of immediate interest in many applied fields, e.g. when investigating whether more racially diverse university programs improve educational outcomes (Antonio et al., 2004) or whether increased income equality boosts economic output (Shin, 2012). The prevalent approach is to summarize diversity via a one-dimensional summary statistic $G(Z)$ (e.g., the Gini coefficient, generalized entropy or Gini-Simpson index) and analyze the correlation between $G(Z)$ and $Y$. As in Example 1 this type of marginal dependence analysis is potentially misleading in practice. In Table 1 we illustrate this on a specific numeric example, where the true causal effect is positive, but the marginal method leads to a negative effect of diversity (details are provided in Section S4.1 of the supplementary material). To avoid this type of confounding bias, one would again like to adjust for the remaining dependence within $Z$. However, since $G(Z)$ depends on all of $Z$ existing adjustment approaches are not applicable. The compositional diversity effect (with Gini-speed) which we introduce in Section 2.2 appropriately adjusts for the confounding as shown in Table 1. A more in-depth example in which the marginal approach is misleading is given in Section 4.1.1 where we estimate the effect of racial diversity on income on a semi-synthetic dataset based on US census data. A real application is given in Section 4.1.2 where the effect of racial diversity on school grades is analyzed using data from schools in New York. Unlike in Example 1, the problem of not being able to adjust with existing techniques is not solely due to the simplex-constraint and similarly appears for summary statistics of unconstrained data. Our perturbation-based framework can also be extended to those settings, but as the use of this type of analysis is particularly prevalent in compositional data applications we restrict our attention to this setting.
@@
We postpone the formal definitions of perturbations and average perturbation effects to Section 2 but give a brief description here. A perturbation is a function $\psi:\mathcal{D}\subseteq\Delta^{d-1}\times[0,\infty)\rightarrow\Delta^{d-1}$ which for each point $z\in\Delta^{d-1}$ describes a trajectory $\gamma\mapsto\psi(z,\gamma)$ along which the point $z$ should be changed. For example, in Example 1 when considering the first microbe (i.e., $j=1$), we could consider a binary perturbation $\psi:\Delta^{d-1}\times\{0,1\}\rightarrow\Delta^{d-1}$ satisfying for all $z\in\Delta^{d-1}$ that $\psi(z,0)=z$ and $\psi(z,1)=(0,\mathsf{C}(z^{-j}))$, where $\mathsf{C}(\cdot)$ scales the inputs to have sum one. By defining $f:z\mapsto\mathbb{E}[Y\,|\,Z=z]$ as the regression function, we can then consider contrasts of the form $f(\psi(z,1))-f(\psi(z,0))$, which capture the effect of setting $z^{1}=0$ while the other coordinates are proportionally rescaled to account for the change in $z^{1}$. The fact that $Z$ is compositional necessitates that we fully specify what happens to all of $Z$ rather than only modifying $Z^{j}$ as one would do in the usual unconstrained approach outlined in Example 1. By marginalizing over $Z$ the perturbation-based approach leads to the average effect
$\mathbb{E}[f(\psi(Z,1))-f(\psi(Z,0))].$
In order to make the effect measure agnostic to the proportion of $Z$s that already have $Z^{j}=0$ in the population, we instead propose to consider the effect measure
@@
$\lambda_{\psi}$ measures the expected change in $Y$ when setting $Z^{j}=0$ for an individual that does not already have $Z^{j}=0$. In Example 2, an increase in the diversity of $Z$ can be expressed as a perturbation $\psi$ that pushes every point in $\Delta^{d-1}$ to a more diverse point. Formally, using the center of the simplex (i.e., $z_{\mathrm{cen}}=(1/d,\ldots,1/d)\in\Delta^{d-1}$) as the most diverse point, we define for all $z\in\Delta^{d-1}$ and all $\gamma\in[0,\|z_{\mathrm{cen}}-z\|_{1}]$ the perturbation
@@
which quantifies how on average the conditional expectation of $Y$ given $Z$ is affected by small increases in diversity. A key strength of using perturbations to define target parameters is that it makes them easy to communicate, interpret and discuss with practitioners. From our experience, practitioners often have a good understanding of what perturbation they are interested in and formalizing them mathematically as above is straight-forward. Even though the target parameters $\lambda_{\psi}$ and $\tau_{\psi}$ have a clear causal motivation, it is important to stress that they are purely (observational) statistical quantities. To connect them to causal effects requires standard causal assumptions and potentially adjustments to the estimators, which we discuss in Section 2.3.
@@
While the target parameters $\lambda_{\psi}$ and $\tau_{\psi}$ may seem simple at first sight, they can be challenging to estimate directly as they depend on estimating the regression function $z\mapsto\mathbb{E}[Y\,|\,Z=z]$ which, in part due to the simplex structure, is generally a complex functional. We therefore propose an approach that reparametrizes the compositional vector $Z\in\Delta^{d-1}$ via a bijection $\phi$ to a vector $(L,W)\in\mathbb{R}\times\mathcal{W}$ in such a way that the perturbation only changes $L$ in the new parametrization. To make this more concrete, assume that the structural causal model (SCM) (e.g., Pearl, 2009) given in Figure 1 (left) generates the data (the causal terminology is used only to provide a more intuitive introduction). Given a perturbation $\psi$ that we are interested in, we now propose to construct a bijection $\phi:\Delta^{d-1}\rightarrow\mathbb{R}\times\mathcal{W}$ that maps $Z$ to $(L,W)$ such that if we intervene on $Z$ by setting it to $\psi(Z,\gamma)$ this corresponds to an intervention in an SCM over the variables $(Y,L,W)$ on $L$ alone (see Figure 1 (right)). Hence, if we are able to construct such a reparametrization, then we have successfully reduced our problem to a setting in which standard adjustment works, that is, using any existing method to estimate the effect of $L$ on $Y$ that adjusts for $W$ results in an unbiased estimate of the causal effect. Given a perturbation $\psi$, we provide a general recipe to construct the reparametrization $\phi$. We furthermore propose several explicit perturbations (including their corresponding reparametrizations) and suggest efficient estimators of the perturbation-based target parameters using semiparametric theory based on the $(L,W)$-reparametrization.
(a) SCM on $Z$-scale (b) SCM on $(L,W)$-scale
$Y$ $\leftarrow f(Z,\eta)$
$Y$$Z$$\psi$$f$
bijection $\phi$$\phi(Z)=(L,W)$$Z=\phi^{-1}(L,W)$
$Y$$L$$W$$\psi$$h$$m$$h$
$L$ $\leftarrow m(W,\xi)$
$Y$ $\leftarrow h(L,W,\varepsilon)$
Figure 1: (a) Data-generating SCM on the $Z$-scale. $Z\in\Delta^{d-1}$ is a compositional vector, $Y$ is a response variable and $\eta$ is exogenous noise. (b) Reparametrized SCM on the $(L,W)$-scale. $(L,W)=\phi(Z)$ are the reparametrized predictors, $Y$ is the same response as in (a) and $\xi$ and $\varepsilon$ are exogenous noise variables. The reparametrization needs to disentangle the perturbation effect on $Z$ in a way such that intervening on $Z$ by setting it to $\psi(Z,\gamma)$ corresponds to intervening on only $L$ (red double head arrows).
The remainder of the paper is structured as follows. In Section 1.1 we discuss existing methods for compositional data analysis and how our approach contributes to this research. Section 1.2 collects the mathematical notation used throughout this work. Our perturbation-based framework is introduced in Section 2. In Section 2.1, we provide a way to construct interpretable perturbations based on suitably choosing the endpoint, direction and speed of the perturbation. We also provide a general construction for the aforementioned $L$ and $W$ for a wide range of perturbations. We end the section by discussing how to adapt our proposals to general geometries on the simplex including those employed by many current methods in compositional data analysis. In Section 2.2 we highlight and name some of the perturbation effects that we believe are most useful for practitioners and provide some examples of how and when these effects should be applied. We provide a short discussion on how to connect the average perturbation effects to causal quantities in Section 2.3. In Section 3, we discuss how to estimate the parameters of interest based on semiparametric theory using two different methods that represent a trade-off between generality and robustness. Our semiparametric estimators and perturbation effects are then illustrated on both simulated and semi-synthetic data and applied to two real-world applications in Section 4. Finally, we present the theoretical guarantees of our estimates in Section 5.
1.1 Related literature
This paper addresses problems from several different fields of statistics. Our proposed perturbation-based framework is perhaps most closely related to the literature on causal inference, semiparametric inference and targeted learning (Bickel et al., 1993; Tsiatis, 2006; van der Laan and Rose, 2011). Spirtes and Scheines (2004) discuss some of the problems of modelling causal effects when certain variables are deterministic functions of other variables. This discussion was later continued by Woodward (2015) and new causal models have been developed to address this problem (Beckers et al., 2023). We contribute to these developments by providing a novel framework for causal effect estimation when some covariates are compositional (see Section 2.3).
An important aspect of causal effect estimation more generally is the need for an interpretable summary measure of the effect. Historically, such interpretable measures originated from parametric models that were easy to interpret. Much work has since gone into the interpretation of and uncertainty quantification for such parameters when the model is misspecified. Perhaps one of the most impactful such analyses was given by White (1980) who provided an estimation procedure for the coefficients in a linear model that did not require the model to be correctly specified. To what extent parameters based on linear models are useful and interpretable when the model is misspecified remains an active research topic particularly in econometrics (Aronow and Samii, 2016; Graham and de Xavier Pinto, 2022; Słoczyński, 2022). More recently, interpretable parameters based on partially linear models have been proposed and analyzed extensively (Chernozhukov et al., 2018; Vansteelandt and Dukes, 2022). Despite these and similar parameters depending on potentially complicated nuisance functions, it is often possible to estimate them efficiently under relatively mild conditions, as we also do here, by utilizing semiparametric theory (Kennedy, 2023).
Model-based effect estimation with compositional variables has a long history and was perhaps first thoroughly investigated under the name ‘experimental design with mixtures’ (Claringbold, 1955; Cornell, 2002). While parts of the literature focused on the problem of selecting design points $z_{1},\dots,z_{n}\in\Delta^{d-1}$ in optimal ways, a key aspect was also the problem of constructing reasonable regression models in this setting. Scheffé (1958) pointed out that polynomial models are overidentified with compositional data due to the sum-to-one constraint and provided a canonical choice of polynomials. Becker (1968) critiqued the use of polynomials for compositional features and instead provided parametric models with homogenous functions of degree one as building blocks. Cox (1971) argued that the polynomials of Scheffé would often not result in interpretable coefficients but that one should instead define the coefficients differently, e.g. so that the $j$th coefficient in a linear model corresponds to increasing the value of $Z^{j}$ while scaling the remaining components towards a fixed and problem-dependent ‘standard mixture’ $z_{*}\in\Delta^{d-1}$. It is possible to define a model-agnostic version of Cox’s linear effects using our perturbation-based framework (see Proposition S4 in Section S6 of the supplementary material). Draper and John (1977) developed predictive polynomial models based on the reciprocal of the components of $Z$ to cope with cases where the expected response diverges at the boundary. Aitchison and Bacon-Shone (1984) introduced the log-contrast regression model where the linear version corresponds to a linear regression on the log-transformed $Z$ with a sum-to-zero constraint on the coefficients. Log-contrast models have seen extensive use in practice and have since been generalized to high-dimensional $Z$ (Lin et al., 2014; Simpson et al., 2021). One of our proposed feature influence measures, the compositional feature influence with multiplicative speed (see Example 3), corresponds exactly to the log-contrast coefficients when the model is correctly specified (see Proposition S3 in Section S6 of the supplementary material or Huang et al. (2023) where this was first shown). Defining practical predictive models for compositional data remains an active topic of research (Xiong et al., 2015; Fiksel et al., 2022; Li et al., 2023; Bhaduri et al., 2025).
Naturally, our proposed average perturbation effects are also connected to the broader literature on compositional data analysis. It is well-known that many conventional statistical procedures become invalid when applied to compositional data. For example, already Pearson (1897) drew attention to the fact that correlations between individual compositional components can be spurious due to the sum-to-one constraint. The most prevalent class of general methods is based on the seminal paper by Aitchison (1982) which introduces the log-ratio approach. In its most basic form it consists of transforming the simplex data via the additive, centered or isometric log-ratio transforms (Aitchison, 1982, 1986; Egozcue et al., 2003) and then performing the statistical analysis in the transformed space. Our approach is not based on log-ratios, but it is possible to construct perturbations from this perspective (see Section 2.1.3). The idea of transforming compositional data via the log-ratio transform to a simpler sample space has been applied successfully in many applications, however this in principle requires $Z$ to take values in the open simplex. This is known as the zero problem and has sparked a large body of research developing various types of zero-imputation schemes (e.g., Martín-Fernández et al., 2000, 2003, 2015; Palarea-Albaladejo and Martín-Fernández, 2008; Lubbe et al., 2021) that can be used before applying the log-ratio approach. Greenacre et al. (2023) provide an overview of current state-of-the-art compositional data analysis and highlights some of the shortcomings in existing transformation-based approaches which are further highlighted in Scealy and Welsh (2014). While many of these approaches are sensible in specific applications, we believe that it is generally preferable to modify the statistical procedure to fit the data rather than vice versa.
There are other approaches to the analysis of data on the simplex, both in the form of alternative transformations, for example power-transformations (Greenacre, 2009) or spherical transformations (Scealy and Welsh, 2011; Li and Ahn, 2022), or by appealing to more general notions of object-oriented data analysis (Marron and Dryden, 2021). There has also been much recent work on manifold-valued data where compositional data is a special case (Dai and Müller, 2018; Dai et al., 2021; Petersen et al., 2022; Kurisu et al., 2024).

### v4

@@ Abstract
Existing effect measures for compositional features are inadequate for many modern applications, for example, in microbiome research, since they display traits such as high-dimensionality and sparsity that can be poorly modelled with traditional parametric approaches. Further, assessing – in an unbiased way – how summary statistics of a composition (e.g., racial diversity) affect a response variable is not straightforward. We propose a framework based on hypothetical data perturbations which defines interpretable statistical functionals on the compositions themselves, which we call average perturbation effects. These effects naturally account for confounding that biases frequently used marginal dependence analyses. We show how average perturbation effects can be estimated efficiently by deriving a perturbation-dependent reparametrization and applying semiparametric estimation techniques. We analyze the proposed estimators empirically on simulated and semi-synthetic data and demonstrate advantages over existing techniques on data from New York schools and microbiome data.
@@ 1 Introduction
Modelling the relationship between a real-valued response $Y\in\mathbb{R}$ and a composition $Z=(Z^{1},\dots,Z^{d})\in\Delta^{d-1}$ has a long history and has, for example, been considered in the literature on ‘experimental design with mixtures’ (Claringbold, 1955; Cornell, 2002) and ’log-ratio based compositional data analysis’ (Aitchison, 1982). We summarize the historical development of these fields in Section S2.1 of the supplementary material. A key aspect of this research is that it attempts to achieve two separate goals simultaneously: (i) ensuring that the model parameters are interpretable and (ii) providing parametric models that fit to realistic data-generating processes with compositional covariates. More recently, in the context of high-dimensional and sparse data, alternative approaches (Knight et al., 2018; Cammarota et al., 2020; Huang et al., 2023) to solving the second goal have been proposed that apply flexible machine learning methods to model the relationship between $Y$ and $Z$ nonparametrically. However, such approaches are at odds with the first goal of providing interpretable effect measures and cannot be combined with the above-mentioned approaches which heavily rely on parametric models to define such effects.
@@
Our main contribution is to provide a unified framework for defining interpretable effect measures for compositional features which we call average perturbation effects that can be combined with flexible nonparametric estimation procedures. Special cases of these average perturbation effects have interpretations that are similar to some model-based individual feature effect measures, e.g., Cox-directions (Cox, 1971) and log-contrast coefficients (Aitchison and Bacon-Shone, 1984). Average perturbation effects, however, are defined without reference to a specific regression model and therefore do not rely on a correctly specified parametric model. Importantly, this allows us to construct more flexible target quantities, two of which we believe are particularly relevant for applications. First, effects based on binary perturbations are useful when summarizing the effect of a discrete change, e.g. the effect of setting $Z^{j}$ to 00 while rescaling $Z^{-j}:=(Z^{1},\dots,Z^{j-1},Z^{j+1},\dots,Z^{d})$ to the simplex. In Section 4.2 we show empirically that the compositional features with significant effects as measured by these binary perturbations coincide with those deemed important by the standard log-contrast regression procedures used in the microbiome literature. Second, effects based on directional perturbations summarize continuous changes in specific directions. This can involve changing individual components of $Z$, but also allows for more complex changes that are not tied to a specific coordinate of $Z$. For example, we can define the effect of ‘diversifying’ $Z$, i.e. moving closer to the center of $\Delta^{d-1}$. Such diversifying effects have been considered previously (Antonio et al., 2004) but are usually defined as the effect of marginally changing a specific measure of diversity. However, as the effect of diversifying can be confounded by other features of $Z$ such a marginal analysis can be misleading and due to the simplex constraint, there is no immediate way to correct for this confounding when estimating these effects (see Section 4.1). In contrast, our perturbation-based framework can control for the remaining variation in $Z$ leading to unconfounded effect estimates with valid asymptotic inference.
@@
Our proposal is intimately related to the growing literature on targeted learning and semiparametric inference (Chernozhukov et al., 2018; Berk et al., 2021; Vansteelandt and Dukes, 2022). An important aspect of our work is, however, to clarify the need for tailored target estimands in the presence of compositional structure in the covariates. Applying a nonparametric effect measure that is not aware of the sum-to-one constraint of the compositional vector can be just as misleading as naively estimating a non-compositional effect based on a parametric model (see Section 4.2). Beyond defining average perturbation effects, we also provide several algorithm-agnostic procedures to estimate the average perturbation effects. These procedures are essentially well-known semiparametric learning algorithms applied after appropriately transforming the compositional covariates. While these algorithms require us to estimate nuisance functions relating to the distribution of $(Y,Z)$, which are particularly delicate for compositional covariates, they only require these estimates to be accurate and not interpretable. This allows us to apply arbitrary regression methods as long as they are sufficiently predictive.
@@
Average perturbation effects are defined by comparing the change in the expected response of $Y$ under a hypothetical perturbation of the composition $Z$. To motivate this, consider unconstrained covariates $X\in\mathbb{R}^{d}$ instead of the composition $Z\in\Delta^{d-1}$. Suppose that there exists $\beta=(\beta^{1},\dots,\beta^{d})\in\mathbb{R}^{d}$ such that for all $x=(x^{1},\dots,x^{d})\in\mathbb{R}^{d}$, $f(x):=\mathbb{E}[Y\,|\,X=x]=\sum_{j=1}^{d}\beta^{j}x^{j}$, that is, the conditional expectation of $Y$ given $X=x$ is a linear function of $x$. Then, one interpretation of the coefficient $\beta^{j}$ is that it corresponds to the expected rate of change of the response in the $j$th coordinate, or more formally, $\partial_{x_{j}}f(x^{1},\dots,x^{d})=\beta^{j}.$ We can rephrase this interpretation by considering a perturbation $\psi:\mathbb{R}^{d}\times[0,1)\rightarrow\mathbb{R}^{d}$ that additively increases the $j$th coordinate, that is, $\psi(x,\gamma)=(x^{1},\dots,x^{j}+\gamma,\dots,x^{d})$. Then, for any $x\in\mathbb{R}^{d}$, we have
@@
Returning now to the composition $Z$ and redefining for $z=(z^{1},\dots,z^{d})\in\Delta^{d-1}$, $f(z):=\mathbb{E}[Y\,|\,Z=z]$, we see that neither of the above expressions are well-defined. The quantity $f(\psi(z,\gamma))$ is meaningless for any non-zero $\gamma\in\mathbb{R}$, since $f$ is only defined on $\Delta^{d-1}$. We can however use similar ideas by defining tailored perturbations on $\Delta^{d-1}$ to obtain interpretable effect measures. In Section S2.2 of the supplementary material we provide two helpful examples to illustrate how conventional effect measures applied to compositional data can be misleading and how our approach mitigates this.
@@
A perturbation, formally defined in Section 2, is a function $\psi:\mathcal{D}\subseteq\Delta^{d-1}\times[0,\infty)\rightarrow\Delta^{d-1}$ which for each point $z\in\Delta^{d-1}$ describes a trajectory $\gamma\mapsto\psi(z,\gamma)$ along which the point $z$ should be changed. For example, we could consider a binary perturbation $\psi:\Delta^{d-1}\times\{0,1\}\rightarrow\Delta^{d-1}$ satisfying for all $z\in\Delta^{d-1}$ that $\psi(z,0)=z$ and $\psi(z,1)=(0,\mathsf{C}(z^{-j}))$, where $\mathsf{C}(\cdot)$ scales the inputs to have sum one. By defining $f:z\mapsto\mathbb{E}[Y\,|\,Z=z]$ as the regression function, we can then consider contrasts of the form $f(\psi(z,1))-f(\psi(z,0))$, which capture the effect of setting $z^{1}=0$ while the other coordinates are proportionally rescaled to account for the change in $z^{1}$. The fact that $Z$ is compositional necessitates that we fully specify what happens to all of $Z$ rather than only modifying $Z^{j}$ as one would do in the usual unconstrained approach. By marginalizing over $Z$ the perturbation-based approach leads to the average effect $\mathbb{E}[f(\psi(Z,1))-f(\psi(Z,0))].$ In order to make the effect measure agnostic to the proportion of $Z$s that already have $Z^{j}=0$ in the population, we instead propose to consider the effect measure
@@
The quantity $\lambda_{\psi}$ measures the expected change in $Y$ when setting $Z^{j}=0$ for an individual that does not already have $Z^{j}=0$. An increase in the diversity of $Z$ can be expressed as a perturbation $\psi$ that pushes every point in $\Delta^{d-1}$ to a more diverse point. Formally, using the center of the simplex (i.e., $z_{\mathrm{cen}}=(1/d,\ldots,1/d)\in\Delta^{d-1}$) as the most diverse point, we define for all $z\in\Delta^{d-1}$ and all $\gamma\in[0,\|z_{\mathrm{cen}}-z\|_{1}]$ the perturbation
@@
which quantifies how on average the conditional expectation of $Y$ given $Z$ is affected by small increases in diversity. A key strength of using perturbations to define target parameters is that it makes them easy to communicate, interpret and discuss with practitioners. From our experience, practitioners often have a good understanding of what perturbation they are interested in and formalizing them mathematically as above is straight-forward. Even though the target parameters $\lambda_{\psi}$ and $\tau_{\psi}$ have a clear causal motivation, it is important to stress that they are purely (observational) statistical quantities. To connect them to causal effects requires standard causal assumptions and potentially adjustments to the estimators, which we discuss in Section S3 of the supplementary material.
@@
While the target parameters $\lambda_{\psi}$ and $\tau_{\psi}$ may seem simple at first sight, they can be challenging to estimate directly as they depend on estimating the regression function $z\mapsto\mathbb{E}[Y\,|\,Z=z]$ which, in part due to the simplex structure, is generally a complex functional. We therefore propose an approach that reparametrizes the compositional vector $Z\in\Delta^{d-1}$ via a bijection $\phi$ to a vector $(L,W)\in\mathbb{R}\times\mathcal{W}$ in such a way that the perturbation only changes $L$ in the new parametrization. Hence, if we are able to construct such a reparametrization, then we have successfully reduced our problem to a setting in which using any existing method to estimate the effect of $L$ on $Y$ that adjusts for $W$ results in an unbiased estimate of the desired effect. Given a perturbation $\psi$, we provide a general recipe to construct the reparametrization $\phi$. We furthermore propose several explicit perturbations (including their corresponding reparametrizations) and suggest efficient estimators of the perturbation-based target parameters using semiparametric theory based on the $(L,W)$-reparametrization.
The paper is structured as follows. Our perturbation-based framework is formally introduced and developed in Section 2. We describe how to construct perturbations in Section 2.1 and provide a table of pre-defined simplex effects along with examples of their use in Section 2.2. In Section 3 we give an overview of the semiparametric techniques used to estimate our proposed effects. In Section 4 we provide a semi-synthetic and two real world applications of our proposal. The supplementary material contains technical details on theory and algorithms along with additional examples and experiments.